OpenStudy (clara1223):

find the limit as x approaches 0 of (sin(4x))/(x(cos(x))) a) -1 b) 0 c) 4 d) 1/4 e) does not exist

OpenStudy (anonymous):

have you tried graphing the equation and trying to solve it graphically yet?

OpenStudy (clara1223):

no, not yet.

OpenStudy (anonymous):

i would suggest that as the answer may become apparent to you when you graph it

zepdrix (zepdrix):

\[\large\rm \lim_{x\to 0}\frac{\sin(4x)}{x \cos x}=\quad \lim_{x\to 0}\frac{\sin(4x)}{x}\cdot \lim_{x\to 0}\frac{1}{\cos x}\]Maybe this can help get us started. When I break them up this way, that first limit looks an awful lot like an identity doesn't it?

OpenStudy (clara1223):

graphically it says that x is undefined at 0 but I need to prove it on paper

OpenStudy (clara1223):

@zepdrix then I can pull a 4 out of the first limit and the answer is 4, I'm not sure how to get the second limit though

zepdrix (zepdrix):

This is how you should think about limits in your brain: Step 1: Plug the limit value directly into the function. Step 2: If there is a problem, back up, and do some algebra. Step 3: Plug the limit directly in and check again.

zepdrix (zepdrix):

So for the second limit, do step 1, .... and you're done with it.

OpenStudy (clara1223):

ok 1/cos(0) is 1 so the overall answer is 4 correct?

zepdrix (zepdrix):

yay good job \c:/